# A physics-informed neural network for modeling fracture without gradient damage: formulation, application, and assessment

### v1

@@ Abstract
Accurate computational modeling of damage and fracture remains a central challenge in solid mechanics. The finite element method (FEM) is widely used for numerical modeling of fracture problems; however, classical damage models without gradient regularization yield mesh-dependent and usually inaccurate predictions. The use of gradient damage with FEM improves numerical robustness but introduces significant mathematical and numerical implementation complexities. Physics-informed neural networks (PINNs) can encode the governing partial differential equations, boundary conditions, and constitutive models into the loss functions, offering a new method for fracture modeling. Prior applications of PINNs have been limited to small-strain problems and have incorporated gradient damage formulation without a critical evaluation of its necessity. Since PINNs in their basic form are meshless, this work presents a PINN framework for modeling fracture in elastomers undergoing large deformation without the gradient damage formulation. The PINN implementation here does not require training data and utilizes the collocation method to formulate physics-informed loss functions. We have validated the PINN’s predictions for various defect configurations using benchmark solutions obtained from FEM with gradient damage formulation. The crack paths obtained using the PINN are approximately insensitive to the collocation point distribution. This study offers new insights into the feasibility of using PINNs without gradient damage and suggests a simplified and efficient computational modeling strategy for fracture problems. The performance of the PINN has been evaluated through systematic variations in key neural network parameters to provide an assessment and guidance for future applications. The results provide motivation for extending PINN-based approaches to a broader class of materials and damage models in mechanics.
@@
The ability to computationally model and predict damage and fracture of materials and structures is essential for enhancing scientific understanding of failures and enabling robust real-world engineering solutions. The finite element method (FEM) is extensively used for numerical modeling of fracture for a variety of problems (Keyak et al., 2001, Wangen, 2011, Oh et al., 2011). A straightforward way to model damage and fracture is using a local damage field variable in FEM without incorporating non-local gradient damage terms. In FEM, the stiffness of an element is reduced to zero when the damage field variable at an integration point reaches a critical value (fully damaged value at the local point). Solutions from FEM without gradient damage have been shown to exhibit significant mesh-dependency. The predicted crack propagation path can be nonphysical and inaccurate if the finite element mesh is not aligned with the actual path, which can be difficult to guess a priori (Song et al., 2008, Pelfrene et al., 2016). The nonlocal gradient damage term in mesh-based numerical modeling is needed to regularize for mesh-independent solutions. Computational models and methods that can predict crack nucleation, propagation, and coalescence in arbitrary, complex geometries without requiring the mesh to be aligned with a priori unknown crack paths are necessary for solving a general class of problems. For numerical reasons, FEM with gradient damage considers a “smeared crack approach”, in which the mathematical framework for damage depends on a damage field variable along with its gradient . A length scale parameter is also used over which varies between zero and one (Miehe et al., 2010). FEM with gradient damage can provide mesh-independent results, provided that the typical element size is less than 0.2 (Narayan and Anand, 2019). Numerous works in the literature have used FEM with gradient damage to model damage initiation, growth, and complete failure in a broad range of materials (Narayan and Anand, 2021, Kuhl et al., 2000, Konica and Sain, 2021, Yang et al., 2024a, Hu et al., 2020, Kumar and Lopez-Pamies, 2020, Konale and Srivastava, 2025). However, the mathematical and numerical implementation complexity of FEM with gradient damage is significantly higher compared to that of the simple FEM without gradient damage. An additional degree of freedom has to be considered at each point in FEM with gradient damage, and the corresponding partial differential equation for damage evolution has to be solved. This involves spatial gradient calculations and the development of complex user-defined elements and user subroutines for implementation within FEM programs (Narayan and Anand, 2019, Konica and Sain, 2021).
@@
Neural networks have been increasingly explored as computational methods to solve engineering mechanics problems previously difficult to analyze (Niu and Srivastava, 2022b, a). Physics-informed neural networks (PINNs) have emerged in the recent literature as a numerical method for solving mechanics problems (Haghighat et al., 2021, Zhang et al., 2022, Bai et al., 2023, Jin et al., 2023, Niu et al., 2023, Hu et al., 2024, Fuhg and Bouklas, 2022, Abueidda et al., 2023, Wang and Yin, 2024, Zhang et al., 2020). Although PINNs are relatively new compared to FEM (Grossmann et al., 2024), they have certain advantages which include their open-source nature and relative ease of implementation (Rezaei et al., 2022). PINNs can solve partial differential equations (PDEs) by encoding the mathematical information explicitly into the loss functions without a need for training data. Some recent works have focused on the application of PINNs to damage and fracture problems in the small-strain regime (Goswami et al., 2022, 2020, Zheng et al., 2022, Ghaffari Motlagh et al., 2023, Manav et al., 2024, Wang et al., 2024, Hu et al., 2024, Kiyani et al., 2024). These works directly adopt the relatively complex mathematical formulation of the gradient damage method to develop the physics-based loss functions. Loss functions in the form of variational energy are more common than separate loss functions that individually enforce the underlying partial differential equations and boundary conditions. Variational energy-based loss functions require numerical integration over the domain to evaluate energy integrals. Variational energy-based PINNs have been developed in (Goswami et al., 2020, Zheng et al., 2022, Manav et al., 2024), which solve the problem incrementally using transfer learning and without training data. Goswami and co-workers (Goswami et al., 2022) proposed a variational form of DeepONet (V-DeepONet), which is trained by imposing the governing equations in a variational form with labeled data. The V-DeepONet, once trained, can rapidly predict the global solution for any initial crack configuration and loading on that domain. Improvements in the V-DeepONet have been presented in (Kiyani et al., 2024). Yi and co-workers proposed a mechanics-informed, model-free symbolic regression framework for solving small-strain fracture problems.
@@
The PINN and physics-informed machine learning-based damage and fracture modeling work in the literature is limited to small strains, and these PINNs directly incorporate the gradient damage formulation. Further, due to the meshless nature of PINNs (Yang et al., 2024b), the gradient damage formulation utilized primarily for numerical reasons (to eliminate mesh dependency in FEM), is not necessary. The main contributions of this work are:
@@
The PINN as a fracture modeling method is evaluated through systematic variations in key parameters. The PINN crack path predictions are approximately insensitive to the distribution of collocation points. We also show that relatively computationally inexpensive one-time step PINN solutions provide reasonably accurate crack paths.
Recently large deformation fracture of elastomers was accurately modeled in (Konale and Srivastava, 2025) using the gradient damage framework with FEM (Konale and Srivastava, 2025). We use the critical free energy-based damage initiation criteria proposed in (Konale and Srivastava, 2025) for our PINN model. In the presence of multiple defects, crack path prediction is complex. Accurate crack path prediction is important for structural applications. We apply our PINN without gradient damage to predict the crack paths for various defect configurations. The plan of the rest of the paper is as follows. The constitutive equations for the modeling of large deformation fracture of elastomers without gradient damage are presented in Section 2, and the corresponding PINN formulation is presented in Section 3. The results of the PINN predictions are validated in Section 4 with a variety of fracture problems using various defect configurations. In Section 5, we provide an assessment of the PINN’s performance for various neural network parameter choices and conclude the paper in Section 6.
@@ 2 Modeling framework and constitutive equations
where , are the ground state shear and bulk modulus, respectively. , , are the distortional part of C, the first invariant of and determinant of F, respectively. The approach discussed in this section should apply to any functional form (hyperelastic model) of choice for . A damage variable is considered to model damage, which can take the values of either 0 (undamaged) or 1 (fully damaged). This is equivalent to instantaneous damage growth upon initiation at a material point experimentally observed in elastomers (Benvidi et al., 2019, Lee et al., 2024). A material point is considered fully damaged ( = 1) if at a material point and any deformation step exceeds a critical value , i.e.,
@@
This split accounts for damage growth being driven only by distortional and tensile dilatational deformation (Amor et al., 2009, Hesammokri et al., 2023). To ensure irreversibility of damage, it is required that
with being an infinitesimal increment in . The first Piola-Kirchoff stress P and Cauchy stress T tensors are then given as
with denoting the deviatoric part of B.
@@ 3.1 Neural network architecture
We now present the procedure for constructing the PINN to evaluate fracture solutions in arbitrary geometries subjected to large deformations. The PINN has an input layer, one or multiple hidden layers, and an output layer. Two-dimensional problems are solved using the plane strain assumption. The problems are cast in a Lagrangian framework with denoting the coordinates of a material point in the reference configuration. A mixed formulation is used wherein the neural network approximates the first Piola-Kirchoff stress field along with the displacement field . The mixed formulation avoids the increased computational cost associated with higher-order partial derivative computations in the displacement-only formulation, along with accuracy and convergence issues (Niu et al., 2023). In addition, the undamaged free energy , considered as a state variable, is also approximated by the PINN. All trainable PINN parameters, including the weights and biases in the neurons of the network, are denoted by the vector . The PINN formulation can then be written as
with denoting direct outputs of the neural network. Specifically, , . Hence, there are a total of 7 outputs for a given X and . Damage in this modeling framework is supposed to take values of either 0 or 1 as seen in equation (2.3). We evaluate the approximate field using and equations (2.3), (2.5) as
This also simplifies the neural network by reducing the number of outputs.
@@ 3.1.1 Loss function construction
denotes the problem’s referential domain. , are the referential body force per unit volume and referential density, respectively. u is the unknown displacement field and is the prescribed displacement field on part of the boundary . P is the unknown first Piola-Kirchoff stress field, is the outward unit normal vector to the referential configuration surface and is the prescribed traction field on the part of the boundary . The PINN consists of a physics-informed loss function encoding the discrete version of the underlying boundary value problem, in this case, equation (3.3). is discretized using points located at . Similarly, and are discretized using points () and points (), respectively. The time domain is discretized into time steps (). Neglecting body forces, the balances of linear and angular momentum in equation (3.3) can be collocated at a given time instant as
@@
The first Piola-Kirchoff stress and undamaged free energy fields are related to the displacement field through constitutive relations in equations (2.2), (2.3), (2.4), (2.5), and (2.6). Hence, the constitutive first Piola-Kirchhoff stress and undamaged free energy fields are defined as
@@
We train the same neural network for each time step without re-initialization, i.e., the trained network parameters for a time step are used as the initial values during training for the next time step. The expanded forms of the individual loss functions that penalize errors in the collocation equations in the least-squares sense can be written as
with subscripts denoting terms associated with stress, undamaged free energy, displacement boundary condition, and traction boundary conditions, respectively. from equation (3.2) after training for every time step is stored in a global variable for use in the next time step’s training calculations. The trained parameters are stored at desired intervals for visualizing the solution’s temporal evolution. The approximate PINN solution for the damage problem at a referential point X and time instant can then be evaluated as
@@ 4 Fracture predictions for various defect configurations
We apply the PINN to model fracture in three plane-strain geometries that capture a variety of defect (notches and holes) configurations. The chosen geometries probe different phenomena and are used in the literature to validate new numerical methods for modeling fracture. The units and material parameters are specified in a dimensionless manner, as is standard in PINN applications (Cai et al., 2021, Cuomo et al., 2022). The geometries are stretched uniaxially at a speed of 2 and have dimensions of . The values for , , and are taken as 0.77, 1.66, and 3, respectively. Note that in equation (3.14) is obtained using , equation (2.6) and the standard relation between P and T which encodes the balance of angular momentum as
@@
The PINN is implemented in the PyTorch framework and has three hidden layers, each with 30 neurons. The optimization algorithm used is ‘Adam’, with an initial learning rate of 0.0001. Using an exponential decay scheduler, the learning rate is gradually decreased during training. The activation function is the ‘Tanh’ function, except for the last layer, where the ‘Softplus’ activation function is used. This configuration is shown to be optimal for capturing sharp interfaces in the computational domain (Niu et al., 2023), here between the undamaged () and completely damaged () regions. Fourier random feature mapping is performed on the neural network input X = as
with B being the Fourier random feature matrix with dimensions x2. Entries in B are sampled from a Gaussian distribution. is then used as an input to the neural network instead of X. This input mapping has been shown to help capture the high-frequency features in PINN solutions (Niu et al., 2023), which are likely to exist here due to the sharp interfaces between undamaged and completely damaged regions. The training is performed with GPU acceleration on the computation cluster provided by the Center for Computation and Visualization (CCV). It is important to note that, due to hyperelasticity, training can be started at any time before damage initiation. This helps reduce the computation time.
The finite element software ABAQUS/Explicit is used to implement FEM with gradient damage to obtain the baseline solutions for the PINN validation. ABAQUS has been used extensively to successfully model a broad variety of complex mechanics problems (Vaishakh et al., 2024, Konale et al., 2023, Bai et al., 2021, Zhong and Srivastava, 2021, Srivastava et al., 2011). The FEM with gradient damage implementation procedure using a combination of a user-defined material subroutine (VUMAT) for the deformation problem and a user-defined element (VUEL) for the damage problem described in (Konale and Srivastava, 2025) is followed here. The gradient damage method’s mathematical formulation is summarized as
@@
The number of elements used in the finite element meshes for each plate geometry is sufficiently large () to ensure good accuracy of the baseline FEM with gradient damage solutions. Specifically, the finite element meshes used give results reasonably close to the converged solutions while not drastically increasing the computation time. The damage governing equations in the PINN [equations (2.3), (2.5)], and the gradient damage method [equation (4.4), (4.5)] are fundamentally different. Hence, the PINN fracture solutions are validated through comparison of the crack path with the baseline solutions. Point-wise and global errors are evaluated in the deformation stage before fracture. The normalized errors for displacement, Cauchy stress, and undamaged free energy are defined as
where the subscript ref denotes quantities obtained from the baseline FEM with gradient damage solutions. FEM without gradient damage solutions for different mesh structures were also obtained using ABAQUS/Explicit to reemphasize their mesh-dependency. For this, the deformation-only model implemented using a VUMAT was used along with the in-built element deletion in ABAQUS. The solutions of the PINN after training on the grid were evaluated on a testing grid consisting of points to test its extrapolation capabilities. For error evaluation and crack path comparison, nodal data exported from ABAQUS were interpolated on the testing grid.

### v4

@@ Abstract
Accurate computational modeling of damage and fracture remains a central challenge in solid mechanics. The finite element method (FEM) is widely used for numerical modeling of fracture problems; however, classical damage models without gradient regularization yield mesh-dependent and usually inaccurate predictions. The use of gradient damage with FEM improves numerical robustness but introduces significant mathematical and numerical implementation complexities. Physics-informed neural networks (PINNs) can encode the governing partial differential equations, boundary conditions, and constitutive models into the loss functions, offering a new method for fracture modeling. Prior applications of PINNs have been limited to small-strain problems and have incorporated gradient damage formulation without a critical evaluation of its necessity. Since PINNs in their basic form are meshless, this work presents a PINN framework for modeling fracture in elastomers undergoing large deformation without the gradient damage formulation. The PINN implementation here does not require training data and utilizes the collocation method to formulate physics-informed loss functions. We have validated the PINN’s predictions for various defect configurations using benchmark solutions obtained from FEM with gradient damage formulation. The crack paths obtained using the PINN are approximately insensitive to the collocation point distribution. This study offers new insights into the feasibility of using PINNs without gradient damage and suggests a simplified and efficient computational modeling strategy for fracture problems. The performance of the PINN has been evaluated through systematic variations in key neural network parameters to provide an assessment and guidance for future applications. The results indicate that in elastomer-like materials, fracture progression and path evolution are primarily governed by the damage initiation criterion. This study provides motivation for extending PINN-based approaches to a broader class of materials and damage models in mechanics.
@@
The ability to computationally model and predict damage and fracture of materials and structures is essential for enhancing scientific understanding of failures and enabling robust real-world engineering solutions. The finite element method (FEM) is extensively used for numerical modeling of fracture for a variety of problems (FEMfracture1, FEMfracture2, FEMfracture3). A straightforward way to model damage and fracture is by using a local damage field variable in FEM without incorporating non-local gradient damage terms. In FEM, the stiffness of an element is reduced to zero when the damage field variable at an integration point reaches a critical value (fully damaged value at the local point). Solutions from FEM without gradient damage have been shown to exhibit significant mesh-dependency. The predicted crack propagation path can be nonphysical and inaccurate if the finite element mesh is not aligned with the actual path, which can be difficult to guess a priori (Elementdellim1, Elementdel1). The nonlocal gradient damage term in mesh-based numerical modeling is needed to regularize for mesh-independent solutions. Computational models and methods that can predict crack nucleation, propagation, and coalescence in arbitrary, complex geometries without requiring the mesh to be aligned with a priori unknown crack paths are necessary for solving a general class of problems. For numerical reasons, FEM with gradient damage considers a “smeared crack approach”, in which the mathematical framework for damage depends on a damage field variable along with its gradient . A length scale parameter is also used over which varies between zero and one (Miehefirstpaper). FEM with gradient damage can provide mesh-independent results, provided that the typical element size is less than 0.2 (Anandquasibrittle). Numerous works in the literature have used FEM with gradient damage to model damage initiation, growth, and complete failure in a broad range of materials (Anand(2021)flowrule, KUHL2000, Shabnamphasefield, Guocytoskeletalpf, Dolbowpf2, KUMAR1, KONALEMoM). However, the mathematical and numerical implementation complexity of FEM with gradient damage is significantly higher compared to that of the simple FEM without gradient damage. An additional degree of freedom has to be considered at each point in FEM with gradient damage, and the corresponding partial differential equation for damage evolution has to be solved. This involves spatial gradient calculations and the development of complex user-defined elements and user subroutines for implementation within FEM programs (Anandquasibrittle, Shabnamphasefield).
Neural networks have been increasingly explored as computational methods to solve engineering mechanics problems previously difficult to analyze (Niu2022UltrasoundNetwork, Niu2022SimulationMeasurements). Physics-informed neural networks (PINNs) have emerged in the recent literature as a numerical method for solving mechanics problems (PINNsoldmech1, PINNsoldmech2, PINNsoldmech3, PINNsoldmech4, SijunplasticityPINN, PINNfracture8, PINNhyperelasticity1, PINNhyperelasticity2, PINNhyperelasticity4, PINNhyperelasticity5, Bai2025, DONG2025, GUO2025). Although PINNs are relatively new compared to FEM (PINNvsFEM1), they have certain advantages, which include their open-source nature and relative ease of implementation (PINNvsFEM2). PINNs can solve partial differential equations (PDEs) by encoding the mathematical information explicitly into the loss functions without a need for training data. The application of PINNs for mechanics and their advantages and disadvantages with respect to the conventional numerical methods (e.g., FEM) are summarized in Table 1.
@@
Some recent works have focused on the application of PINNs to damage and fracture problems in the small-strain regime (PINNfracture1, PINNfracture2, PINNfracture3, PINNfracture4, PINNfracture5, PINNfracture7, PINNfracture8, PINNfracture9). These works directly adopt the relatively complex mathematical formulation of the gradient damage method to develop the physics-based loss functions. Loss functions in the form of variational energy are more common than separate loss functions that individually enforce the underlying partial differential equations and boundary conditions. Variational energy-based loss functions require numerical integration over the domain to evaluate energy integrals. Variational energy-based PINNs have been developed in (PINNfracture2, PINNfracture3, PINNfracture5), which solve the problem incrementally using transfer learning and without training data. Goswami and co-workers (PINNfracture1) proposed a variational form of DeepONet (V-DeepONet), which is trained by imposing the governing equations in a variational form with labeled data. The V-DeepONet, once trained, can rapidly predict the global solution for any initial crack configuration and loading on that domain. Improvements in the V-DeepONet have been presented in (PINNfracture9). Yi and co-workers proposed a mechanics-informed, model-free symbolic regression framework for solving small-strain fracture problems.
@@
The PINN and physics-informed machine learning-based damage and fracture modeling work in the literature is limited to small strains, and these PINNs directly incorporate the gradient damage formulation. Further, due to the meshless nature of PINNs (PINNmeshless1), the gradient damage formulation utilized primarily for numerical reasons (to eliminate mesh dependency in FEM) is not necessary.
| PINNs for Mechanics |
| • Spatial coordinates, time, PDE parameters, etc., as inputs to the NN. |
| • Displacements, stress tensor components, field variables, PDE parameters, etc., as the NN outputs. |
| • Loss function divided into two parts: data loss (IC/BCs, and any experimental data) and residual/governing equation loss (computed using the strong or weak formulation of PDEs). |
| • Loss function minimization (e.g., MSE, MAE, Sobolev norm) for obtaining the optimal parameters of the NN. The loss is first computed point-wise (for strong formulation) over identical and independently distributed residual (collocation) points as well as IC/BC points. |
| Advantages | Disadvantages |
| Relative ease of implementation from scratch and open-source nature compared to the conventional numerical methods. | Accuracy of the solution depends on various hyperparameters of the neural network. Lack of guidance for a general class of mechanics problems. Rigorous approaches such as grid search can be used for parameter tuning. |
| Mesh-free nature and automatic differentiation (AD) capabilities. AD computes exact gradients as the activation functions have analytical gradient forms. | Forward problems without training data need significantly larger computation time than conventional numerical methods. |
| Non-linear activation functions (e.g., tanh with guaranteed differentiability) are more expressive than conventional method basis functions. | Non-convex, non-linear, and multi-objective nature of loss functions often results in slow convergence. |
| Combination with data-driven learning is promising. Training can be achieved for sparse and high-dimensional data. | “Black box” nature lacks interpretability but can be used as a surrogate model for “Gray box” learning, such as SR/SINDy. |
| Handle ill-posed inverse problems better than conventional methods due to the incorporation of physics as a regularizer - mimetic to Tikhonov regularization. | Application to multi-scale and multi-physics problems is challenging. |
| Can recover missing physics or constitutive relations. Unlike conventional methods, IC/BCs can be partially described. | Detailed validations for a broad class of problems are still in their infancy. |
(IC - initial conditions, BC - boundary conditions, MSE - mean squared error, MAE - mean absolute error, SR - symbolic regression, SINDy - sparse identification of nonlinear dynamics)
The main contributions of this work are:
@@
The PINN as a fracture modeling method is evaluated through systematic variations in key parameters. The PINN crack path predictions are approximately insensitive to the distribution of collocation points. We also show that relatively computationally inexpensive one-time step PINN solutions provide reasonably accurate crack paths for rate-independent materials under monotonic loading.
Recently, large deformation fracture of elastomers was accurately modeled in (KONALEMoM) using the gradient damage framework with FEM (KONALEMoM). We use the critical free energy density-based damage initiation criteria proposed in (KONALEMoM) for our PINN model. In the presence of multiple defects, crack path prediction is complex. Accurate crack path prediction is important for structural applications. We apply our PINN without gradient damage to predict the crack paths for various defect configurations. The plan of the rest of the paper is as follows. The constitutive equations for the modeling of large deformation fracture of elastomers without gradient damage are presented in Section 2, and the corresponding PINN formulation is presented in Section 3. The results of the PINN predictions are validated in Section 4 with a variety of fracture problems using various defect configurations. In Section 5, we provide an assessment of the PINN’s performance for various neural network parameter choices and conclude the paper in Section 6.
@@ 2 Modeling framework and constitutive equations
where , are the ground state shear and bulk modulus, respectively. , , are the distortional part of C, the first invariant of and determinant of F, respectively. The approach discussed in this section should apply to any functional form (hyperelastic model) of choice for . A damage variable is considered to model damage, which can take the values of either 0 (undamaged) or 1 (fully damaged). This is equivalent to instantaneous damage growth upon initiation at a material point experimentally observed in elastomers (EPDMrubber, Parabolicshape2). A material point is considered fully damaged ( = 1) if at a material point at any deformation step exceeds a critical value , i.e.,
@@
This split accounts for damage growth being driven only by distortional and tensile dilatational deformation (HDsplit1, HDsplit2). To ensure irreversibility of damage, it is required that
with being an infinitesimal increment in . The first Piola-Kirchhoff stress P and Cauchy stress T tensors are then given as
with denoting the deviatoric part of .
@@ 3.1 Neural network architecture
We now present the procedure for constructing the PINN to evaluate fracture solutions in arbitrary geometries subjected to large deformations. The PINN has an input layer, one or multiple hidden layers, and an output layer. Two-dimensional problems are solved using the plane strain assumption. The problems are cast in a Lagrangian framework with denoting the coordinates of a material point in the reference configuration. A mixed formulation is used wherein the neural network approximates the first Piola-Kirchhoff stress field along with the displacement field . The mixed formulation avoids the increased computational cost associated with higher-order partial derivative computations in the displacement-only formulation, along with accuracy and convergence issues (SijunplasticityPINN). In addition, the undamaged free energy density , considered as a state variable, is also approximated by the PINN. All trainable PINN parameters, including the weights and biases in the neurons of the network, are denoted by the vector . The PINN formulation can then be written as
with denoting direct outputs of the neural network. Specifically, , . Hence, there are a total of 7 outputs for a given X and . Damage in this modeling framework can take values of either 0 or 1 as seen through equation (2.3). We evaluate the approximate field during training using [using equation (2.4)] and equations (2.3), (2.5) as
@@ 3.1.1 Loss function construction
denotes the problem’s referential domain. , are the referential body force per unit volume and referential density, respectively. u is the unknown displacement field and is the prescribed displacement field on part of the boundary . P is the unknown first Piola-Kirchhoff stress field, is the outward unit normal vector to the referential configuration surface and is the prescribed traction field on the part of the boundary . The PINN consists of a physics-informed loss function encoding the discrete version of the underlying boundary value problem, in this case, equation (3.3). is discretized using points located at . Similarly, and are discretized using points () and points (), respectively. The time domain is discretized into time steps (). Neglecting body forces, the balances of linear and angular momentum in equation (3.3) can be collocated at a given time instant as
@@
The first Piola-Kirchhoff stress and undamaged free energy density fields are related to the displacement field through constitutive relations in equations (2.2), (2.3), (2.4), (2.5), and (2.6). Hence, the constitutive first Piola-Kirchhoff stress and undamaged free energy density fields are defined as
@@
The expanded forms of the individual loss functions that penalize errors in the collocation equations in the least-squares sense can be written as
with subscripts denoting terms associated with stress, undamaged free energy density, displacement boundary condition, and traction boundary conditions, respectively. from equation (3.2) after training for every time step is stored in a global variable for use in the next time step’s training calculations. The trained parameters are stored at desired intervals for visualizing the solution’s temporal evolution. The approximate PINN solution for the damage problem at a referential point X and time instant can then be evaluated for post-processing as
@@
For the fracture problems considered in this work, the inertial terms have a negligible effect, and they are neglected in equation (3.11) used for the PINN loss function formulation. In the baseline FEM solutions, we have incorporated the inertial terms and show that this simplification has negligible effects. The same neural network is trained for each time step without re-initialization, i.e., the trained network parameters for a time step are used as the initial values during training for the next time step.
@@ 4 Fracture predictions for various defect configurations
We apply the PINN to model fracture in three plane-strain geometries that capture a variety of defect (notches and holes) configurations. The chosen geometries probe different phenomena and are used in the literature to validate new numerical methods for modeling fracture. The units and material parameters are specified in a dimensionless manner, as is standard in PINN applications (PINNdimensionless1, PINNdimensionless2). The geometries are stretched uniaxially at a speed of 2 and have dimensions of . The values for , , and are taken as 0.77, 1.66, and 3, respectively. Note that in equation (3.14) is obtained using , equation (2.6) and the standard relation between P and T which encodes the balance of angular momentum as
@@
The PINN is implemented in the PyTorch framework and has three hidden layers, each with 30 neurons. The loss weights , , are set equal to 1. A discussion on this choice can be found in the Appendix. The optimization algorithm used is ‘Adam’, with an initial learning rate of 0.0001. Using an exponential decay scheduler, the learning rate is gradually decreased during training. The activation function is the ‘Tanh’ function, except for the last layer, where the ‘Softplus’ activation function is used. This configuration is shown to be optimal for capturing sharp interfaces in the computational domain (SijunplasticityPINN), here between the undamaged () and completely damaged () regions. Fourier random feature mapping is performed on the neural network input X = as
with B being the Fourier random feature matrix with dimensions 2. Entries in B are sampled from a Gaussian distribution. is then used as an input to the neural network instead of X. This input mapping has been shown to help capture the high-frequency features in PINN solutions (SijunplasticityPINN), which are likely to exist here due to the sharp interfaces between undamaged and completely damaged regions. The training is performed with GPU acceleration on the computation cluster provided by the Center for Computation and Visualization (CCV). It is important to note that, due to hyperelasticity, training can be started at any time before damage initiation. This helps reduce the computation time.
The finite element software ABAQUS/Explicit is used to implement FEM with gradient damage to obtain the baseline solutions for the PINN validation. ABAQUS has been used extensively to successfully model a broad variety of complex mechanics problems (ABAQUSvaishakh, KonalePBS, Bai2021AEngineering, Zhong2021AGrowth, Srivastava2011STRESSPIPE). The FEM with gradient damage implementation procedure using a combination of a user-defined material subroutine (VUMAT) for the deformation problem and a user-defined element (VUEL) for the damage problem described in (KONALEMoM) is followed here. The gradient damage method’s mathematical formulation is summarized as
@@
The number of elements used in the finite element meshes for each plate geometry is sufficiently large (10,000) to ensure good accuracy of the baseline FEM with gradient damage solutions. Specifically, the finite element meshes used give results reasonably close to the converged solutions while not drastically increasing the computation time. The damage governing equations in the PINN [equations (2.3), (2.5)], and the gradient damage method [equation (4.4), (4.5)] are fundamentally different. Hence, the PINN fracture solutions are validated through comparison of the crack path with the baseline solutions. Point-wise and global errors are evaluated in the deformation stage before fracture. The normalized errors for displacement, Cauchy stress, and undamaged free energy density are defined as
where the superscript ref denotes quantities obtained from the baseline FEM with gradient damage solutions. FEM without gradient damage solutions for different mesh structures were also obtained using ABAQUS/Explicit to reemphasize their mesh-dependency. For this, the deformation-only model implemented using a VUMAT was used along with the in-built element deletion in ABAQUS. The solutions of the PINN after training on the grid were evaluated on a testing grid consisting of points to test its extrapolation capabilities. For error evaluation and crack path comparison, nodal data exported from ABAQUS were interpolated on the testing grid.